Hi. Would you please tell me how long I need to make a pipe container to hold 57 litres of milk if the pipe is 8 inches id.
Thanks
John

Answer by jim_thompson5910(35256)

Did you mean 8 inches in diameter? If so, then the radius is 8/2 = 4, so r = 4.


57 liters = 3478.35 cubic inches ----> V = 3478.35




V = pi*r^2*h

3478.35 = pi*4^2*h

3478.35 = 16pi*h

3478.35/16pi = h

h = 3478.35/16pi

h = 682.972425413379

So the pipe must be roughly 682.972425413379 inches long.